OpenStudy (anonymous):

what is the sum of the measures of the interior angles of a 14-sided polygon

OpenStudy (danjs):

sum of interior angles of any polygon = (n-2)*180 where n = number of sides

OpenStudy (danjs):

do you need that formula proved?

OpenStudy (anonymous):

Yes, please

OpenStudy (danjs):

Do you need to know the answer, or the derivation of the formula for the interior angles of a polygon

OpenStudy (danjs):

sum of the interior angles of a 14 sided polygon = (14-2)*180 = 12*180 = 2160

OpenStudy (danjs):

The Exterior angles will always add up to 360

OpenStudy (anonymous):

i have a question what does it mean the number 2 ? I mean (n-2) ? Can you explain me im doing a study guide

OpenStudy (danjs):

Ok The sum of the interior angles of a triangle is 180 right?

OpenStudy (anonymous):

Yes

OpenStudy (danjs):

so a square(4 sided) can be made into 2 triangles by dividing it in half diagonally. so a square has 2 triangles or 2*180 = 360 degrees total interior angles understand that?

OpenStudy (danjs):



OpenStudy (danjs):

3 triangles have 3*180 degrees as the sum of the intoerior angles right?

OpenStudy (danjs):

so a 5 sided figure has 3*180 total interior angles measure

OpenStudy (danjs):

the same process can be done for 6 7 8 9 or n sided polygons,

OpenStudy (danjs):

the trend is always the number of sides subtract 2 multiplied by 180 (number of sides - 2) x 180

OpenStudy (danjs):

it stems from the fact that you know the sum of the interior angles of a triangle adds to 180 degrees.

OpenStudy (anonymous):

Now makes sense, thank you , and i have more questions can you help me ?

OpenStudy (danjs):

k

OpenStudy (anonymous):

?

OpenStudy (danjs):

hi